OpenStudy (anonymous):

what is the derivative of f(x) 2x / x + 2

OpenStudy (anonymous):

I know its f'g - fg' / g^2

OpenStudy (anonymous):

but that is the derivative of x+ 2

OpenStudy (solomonzelman):

\(\large\color{black}{f(x)= \frac{\LARGE 2x }{\LARGE x+2} }\) like this?

OpenStudy (anonymous):

yeah

OpenStudy (solomonzelman):

then I would re-write it as: \(\large\color{black}{f(x)= (2x)(x+2)^{-1} }\) and use the product rule.

OpenStudy (anonymous):

I have to use the quotient rule because that is what the direction says

OpenStudy (anonymous):

And I like quotient rule more I understand it better

OpenStudy (anonymous):

so far I have 2(x+2) - 2x

OpenStudy (zzr0ck3r):

then for \(f(x) = \frac{g(x)}{h(x)}\) we have \(f'(x)=\frac{g'(x)h(x)-g(x)h'(x)}{h(x)^2}\)

OpenStudy (solomonzelman):

the quotient rule is: \(\large\color{black}{ \frac{\LARGE g(x)f'(x)-f(x)g'(x)}{\LARGE g(x)} }\)

OpenStudy (zzr0ck3r):

but that over (x+2)^2

OpenStudy (zzr0ck3r):

put that over*

OpenStudy (anonymous):

what is the derivative of x + 2 ?

OpenStudy (mertsj):

1

OpenStudy (zzr0ck3r):

you are almost done, you forgot the easy part \(\frac{2(x+2)-2x}{(x+2)^2}=\frac{2x+4-2x}{(x+2)^2}=\frac{4}{(x+2)^2}\)

OpenStudy (anonymous):

how do I simplify 2- 2x / x+ 2

OpenStudy (zzr0ck3r):

\(\huge \uparrow\)

OpenStudy (anonymous):

wait I did not do it that way :(

OpenStudy (anonymous):

I got to the second part but instead of getting 4 there is (x +2) on the top and bottom so I took it off

OpenStudy (anonymous):

well I divided it and got 2- 2x / x+ 2